OpenStudy (anonymous):

On the graph of the equation y = 5x2 - 8x + 11, the vertex is the point.

OpenStudy (anonymous):

y' = 0 y' = 10x-8 0=10x-8 x=8/10 y = 5(4/5)^2-8(4/5)+11

OpenStudy (anonymous):

x = 4/5 y = 39/5 vertex = (4/5, 39/5)